### v1

# xBound: Join Size Lower Bounds

Abstract.
Cloud database vendors invest substantial resources into their query optimizers, and for good reason. Cardinality estimation, a cornerstone of the optimizer, is critical for the selection of efficient query plans, as well as downstream tasks such as resource allocation and query scheduling. Yet, as many practitioners and researchers have noted, it is also the optimizer’s Achilles heel. Prior studies on a number of industrial-strength databases show substantial cardinality estimation errors on all tested systems, with a far greater tendency to underestimate than to overestimate. Unfortunately, cardinality underestimation is more problematic than overestimation, as it misleads the optimizer to choose plans designed for small data, leading to underprovisioned CPU and memory.
While previous work on pessimistic cardinality estimation has proposed provable join size upper bounds, such methods can only correct overestimation, leaving the more harmful problem of underestimation unaddressed. To fill this critical gap, we introduce xBound, the very first framework for deriving provable join size lower bounds. xBound successfully reduces underestimation in real systems: On the JOBlight benchmark, it corrects 17.5% of subexpression underestimates in DuckDB and 8.7% in PostgreSQL, while on a Microsoft enterprise workload, it fixes 36.1% of Fabric Data Warehouse’s underestimates, demonstrating a significant step towards solving this long-standing problem.
1. Introduction
Figure 1. Overview of cardinality bound violations across all join subexpressions in JOBlight, when estimating with DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse.
After five decades of research and development, query optimization remains an “unsolved” problem. As pointed out by both practitioners and researchers, cardinality estimation, one of the crucial pillars of the query optimizer, is its Achilles heel (Lohman, 2014, 2017; Leis et al., 2015, 2025). By estimating the size of intermediate results, cardinality estimation is the primary input to the cost model for choosing the best execution plan. Additionally, it is essential for resource management, such as allocating memory and CPU, and determining the degree of parallelism in distributed databases. A seminal analysis by Leis et al. (Leis et al., 2015) empirically evaluated PostgreSQL (Stonebraker and Rowe, 1986), HyPer (Kemper and Neumann, 2011), and three commercial databases, revealing that all tested database systems routinely produce large cardinality estimation errors and that join size estimation is particularly fragile.
This challenge has spurred a wealth of research on cardinality estimation, spanning sketch-based, sampling-based, to ML-based methods (see recent experimental surveys (Harmouch and Naumann, 2017; Han et al., 2021; Heinrich et al., 2025)). However, no matter how well a method has been shown to perform empirically, there is rarely any guarantee on the correctness of the estimation. In response, the theory community has recently proposed provable bounds for cardinality estimation, most notably the recent LpBound (Zhang et al., 2025). Yet, all these proposals provide only upper bounds. Rather than serving as direct replacements for the cardinality estimator, these upper bounds are best suited for correcting the original optimizer’s overestimates (by capping the estimates). Crucially, however, the complementary problem, safeguarding against underestimates, remains unaddressed.
Motivation. Interestingly, as shown in the same analysis (Leis et al., 2015), cardinality underestimation is far more prevalent than overestimation in practice, and this tendency increases with the number of joins. Moreover, although both misestimations are problematic, we argue that underestimation is generally more dangerous. It tricks the optimizer to choose plans best for small data, such as preferring nested-loops joins over hash- or merge-joins, broadcast joins over shuffle joins, causing catastrophic plan choices and order-of-magnitude performance slowdown. Regarding resource management, underestimation risks runtime resource exhaustion, leading to out-of-memory errors or excessive disk spilling. Conversely, overestimation misleads the optimizer to select conservative plans that are safe for big data, as well as overallocating runtime resources. Although this leads to degraded performance and wasted resources, the consequences are seldom catastrophic.
In this paper, we address the critical problem of safeguarding cardinality underestimation, by introducing xBound, the very first theoretical framework for computing provable join size lower bounds.
Technical Overview. xBound, exploits, similar to LpBound (and sketch-based estimators (Wang et al., 2023; Heddes et al., 2024)), the key fact that the size of a join is the dot-product of the join-key degree vectors $\mathbf{a}$ and $\mathbf{b}$ of two relations $A$ and $B$, respectively. Our approach relies on the following observation: Lower bounds require inequalities that bound from below this very dot product; we refer to these as reverse inequalities. Fortunately, reverse inequalities only require a handful of statistics: $\ell_{\infty}$—the maximum key frequency, $\ell_{2}$—the Euclidean norm of the degree vector, and $\ell_{-\infty}$–the minimum key frequency. On the other hand, they require positive input vectors. We will thus first lower bound the number of non-zero entries in the entry-wise product $\mathbf{a}\odot\mathbf{b}$, enabling their application. This step requires the presence of min/max-values on the corresponding join columns, which are readily available in database catalogs and modern file formats.
To get a feeling how insightful lower and upper bounds can be in a real system, consider Fig. 1. We ran xBound and LpBound on all join subexpressions of JOBlight and compared their results against the estimates produced by DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse (DW). Notably, xBound can correct 14.9% of DuckDB’s estimates and 6.9% of PostgreSQL’s. In comparison, LpBound corrects 8.3% and 0.4% of the estimates in the two systems, respectively. Evidently, this dual-bound approach enables us to both detect and mitigate more misestimates during query optimization.
Contributions. We summarize our contributions in the following:
  • 1.
    We introduce xBound, the first framework for provable, non-trivial join size lower bounds, filling a critical gap in the cardinality estimation literature. In its current form, xBound estimates multi-way joins over the same join key.
  • 2.
    We extend our theoretical framework to obtain lower bounds over filtered table scans, supporting equality and range predicates as well as conjunctions and disjunctions.
  • 3.
    We empirically evaluate the impact of xBound in detecting and correcting cardinality underestimates in DuckDB v1.4, PostgreSQL 18, and Fabric DW, across the JOBlight (Kipf et al., 2019b), JOBjoin (Zhang et al., 2025), and STATS-CEB (Han et al., 2021) benchmarks, and on a subset of a Microsoft enterprise workload, MSSalesJoin (Alotaibi et al., 2024).
Structure. The rest of the paper is structured as follows: We provide the preliminaries needed in §2, including $\ell_{p}$-norms and dot-product inequalities. Then, in §3, we detail xBound’s components. Subsequently, we extend our framework to support lower bounds on filtered table scans in §4. We continue with the evaluation of xBound’s lower bounds in §5, and then with a discussion in §6. Afterwards, we outline related work in §7 and conclude in §8.
2. Preliminaries
Notation. For a given relation $R$, denote by $\texttt{Attrs}(R)$ its set of attributes. Let $|\cdot|$ denote the cardinality of a given set. To denote bounds on the cardinality, we use $|\cdot|^{-}$ and $|\cdot|^{+}$ to mean a lower and upper bound, respectively. We use bold characters for vectors, e.g., $\mathbf{a}$. To access position $k$ in a vector $\mathbf{a}$, we write $\mathbf{a}[k]$. For a set of positions $K=\{k_{1},\ldots,k_{m}\}$, we write $\mathbf{a}[K]$ to denote the vector of values at the corresponding positions. In turn, to index a sequence $d$, we use subscript notation, i.e., $d_{k}$ (this distinction helps with readability in the proofs). For integers $a$ and $b$, we define $a,b]\triangleq\{a,\ldots,b\}$, and, as shorthand, $n]\triangleq[1,n]$.
Supported Query Type. Our framework supports select-project-join queries, as follows:
select [projection-list]
from R_1, ..., R_n
where [inner-join-predicates] and [table-predicates];
Throughout the paper, we will assume that the queries have bag semantics, as it is currently the case for leading database systems. [^1] We leave it as future work to understand the effect of our framework for set semantics. To ensure notation consistency with prior work on join size upper bounds, we closely follow the notation in Ref. (Zhang et al., 2025), i.e., we will use the conjunctive query notation instead of SQL:
$Q(V_{0})=R_{1}(V_{1})\land R_{2}(V_{2})\land\ldots\land R_{n}(V_{n})\land[\text{table-predicates}],$
where each $V_{i}\triangleq\texttt{Attrs}(R_{i})$ is the corresponding set of variables of relation $R_{i}$, and $V_{0}\subseteq V_{1}\cup\ldots\cup V_{n}$ are the group-by attributes. Note that, since we will be working with full conjunctive queries only, it holds that $V_{0}=V_{1}\>\cup\ldots\cup\>V_{n}$, i.e., the query does not have a group-by clause. To this end, let $\texttt{Vars}(Q)\triangleq V_{1}\>\cup\ldots\cup\>V_{n}$ denote the set of query variables, and we will write $Q(\ast)$ as a syntactic sugar for a full conjunctive query, i.e., $V_{0}=\texttt{Vars}(Q)$.
Moreover, query $Q$ is acyclic if its relations $R_{1},\ldots,R_{n}$ can be placed on the nodes of a tree such that, for each variable $X_{i}\in\texttt{Vars}(Q)$, the set of tree nodes containing $X_{i}$ forms a connected component. In addition, $Q$ is called Berge-acyclic if it is acyclic and any two relations share at most one variable. For instance, the three-way joins $J_{1}$ and $J_{2}$ are Berge-acyclic, yet $J_{3}$ is cyclic:
(1) $J_{1}(\ast)$ $=R_{1}(X,Y)\land R_{2}(Y,Z)\land R_{3}(Z,U),$
(2) $J_{2}(\ast)$ $=R_{1}(X,Y)\land R_{2}(X,Z)\land R_{3}(X,U),$
(3) $J_{3}(\ast)$ $=R_{1}(X,Y)\land R_{2}(Y,Z)\land R_{3}(Z,X).$
In this work, we focus solely on full Berge-acyclic conjunctive queries, with the additional restriction that joins are over the same join key. Notably, this is the case for many subexpressions of the well-known benchmarks, such as JOB (Leis et al., 2015; Kipf et al., 2019b) and STATS-CEB (Han et al., 2021).
Degree Sequences & Vectors. The degree sequence from $X$ to $Y$ in relation $R$ is the sequence $\deg(Y|X)\triangleq(d_{1},\ldots,d_{N})$, constructed as follows: Compute the domain of $X$, $\texttt{Dom}(R.X)=\{x_{1},\ldots,x_{N}\}$, and denote by $d_{i}=|\sigma_{X=x_{i}}(\Pi_{XY}(R))|$ the degree, i.e., the frequency, of $x_{i}$, and sort the values in ascending order $d_{1}\leq\ldots\leq d_{n}$; note that, in LpBound, the degrees are in descending order. When $|X|\leq 1$, we call the degree sequence simple, and when $XY=\texttt{Attrs}(R)$, we say the degree sequence is full, and denote it by $\deg_{R}(\ast|X)$; we will skip the subscript when it is clear from the context. As constrained by our supported query type (see the above paragraph), we will be using only simple full degree sequences.
$\ell_{p}$-Norms. Instead of storing the degree sequences themselves, we will be storing simple statistics related to them, namely $\ell_{p}$-norms. The advantage of operating on norms is that we have to store only a few bytes per relation; in addition, $\ell_{p}$-norms capture surprisingly well the skew of the degree sequence. Formally, the $\ell_{p}$-norm of a sequence $d=(d_{1},d_{2},\ldots)$ is $\|d\|_{p}\triangleq(\sum_{i}d_{i}^{p})^{1/p}$, where $p\in(0,\infty]$. Note that as $p$ goes to $\infty$, the value of the norms decreases and converges to $\max_{i}d_{i}$. Finally, notice that the $\ell_{p}$-norms are invariant to the order of the elements in the underlying sequence. In addition, as a syntactic sugar, we introduce $\ell_{-\infty}\triangleq\min_{i}d_{i}$. Usually, this is 1, yet there are relations in the benchmarks where this is larger, e.g., movie_info_idx’s movie_id key-column in the IMDb dataset (Leis et al., 2015).
To denote the $\ell_{p}$-norm of a given column $X$ in relation $R$, we write $\ell_{R,X,p}$. As we cannot always compute the norm values exactly, we will work with lower and upper bounds thereof. To this end, we write $\ell_{p}^{-}$ and $\ell_{p}^{+}$ to denote its lower and upper bound, respectively.
To understand how prior work on join size upper bounds utilizes $\ell_{p}$-norms, we refer the reader to related work (see  §7).
2.1. Rearrangement Inequality
Before we move on to the dot-product reverse inequalities that we will be using, let us consider an inequality that connects dot-products of arbitrarily ordered vectors with dot-products of ordered sequences. This will turn useful for the overall understanding of xBound (and its difference to LpBound).
Inequality 1 (Rearrangement Inequality).
Let $a_{1}\leq\ldots\leq a_{d}$ and $b_{1}\leq\ldots\leq b_{d}$ be two non-decreasing sequences of real numbers. Then, for every permutation $\sigma$ of $[d]$, it holds that
$\sum_{i=1}^{d}a_{i}\,b_{d+1-i}\;\leq\;\sum_{i=1}^{d}a_{i}\,b_{\sigma_{i}}\;\leq\;\sum_{i=1}^{d}a_{i}\,b_{i}.$
In a nutshell, the RHS of the inequality is an upper bound on the dot product, while the LHS is a lower bound. In fact, the RHS is used in SafeBound (Deeds et al., 2023b, a), a pre-runner of LpBound (Zhang et al., 2025) (see  §7 for an exposition). What we (implicitly) leverage in xBound is the LHS, i.e., taking the dot-product between reversely ordered sequences. Yet, similar to LpBound, we will be directly bounding the dot-product by $\ell_{p}$-norms only.
2.2. Dot-Product Reverse Inequalities
As motivated in the introduction, we need the so-called reverse inequalities on the dot product of two degree vectors $\mathbf{a}$ and $\mathbf{b}$. The caveat is that this class of inequalities requires the input vectors to be strictly positive, i.e., $a_{i},b_{i}>0,\forall i\in[d]$. Recall that degree vectors are generally only non-negative, i.e., they contain zeros for non-existing keys in a relation. In  §3.1, we show how we can still apply these inequalities by lower-bounding the number of non-zero entries in both vectors.
We start with a well-known reverse inequality:
Inequality 2 (Pólya–Szegő Inequality (Pólya and Szegő, 1925)).
Let $\mathbf{a}=(a_{1},\ldots,a_{d})$ and $\mathbf{b}=(b_{1},\ldots,b_{d})$ be two positive vectors of real numbers such that
$0<m_{a}\leq a_{i}\leq M_{a}<\infty,\;0<m_{b}\leq b_{i}\leq M_{b}<\infty,\;\forall i\in[d].$
Then
$\sum_{i=1}^{d}a_{i}b_{i}\;\geq\;2\,\frac{\sqrt{\displaystyle\sum_{i=1}^{d}a_{i}^{2}}\;\sqrt{\displaystyle\sum_{i=1}^{d}b_{i}^{2}}}{\sqrt{\dfrac{M_{a}M_{b}}{m_{a}m_{b}}}+\sqrt{\dfrac{m_{a}m_{b}}{M_{a}M_{b}}}}.$
Let us inspect this inequality. To lower-bound the dot product $\mathbf{a}\cdot\mathbf{b}$, it requires the following vector norm statistics:
  • $\ell_{2}$: $\|a\|_{2}$ and $\|b\|_{2}$,
  • $\ell_{\infty}$: $M_{a}\triangleq\max_{i}a_{i}$ and $M_{b}\triangleq\max_{i}b_{i}$,
  • $\ell_{-\infty}$: $m_{a}\triangleq\min_{i}a_{i}$ and $m_{b}\triangleq\min_{i}b_{i}$.
As we will estimate the cardinalities of $k$-way joins, we also need the generalization of the above inequality for $k$ vectors. Fortunately, there is one. This kind of inequalities forms the interesting class of reverse Hölder’s inequalities. As a backbone, we will be using the (currently) tightest inequality in the literature, which indeed generalizes Ineq. (2). Hence, it infers the limitation of relying on $\ell_{2}$ only. We discuss possible extensions in  §6.
Figure 2. Obtaining a lower bound on the join size, assuming all keys join, by multiplying cross-wise the values of the (non-decreasing) degree sequences. This follows from the rearrangement inequality, Ineq. (1).
Inequality 3 (Generalized Reversed Hölder’s Inequality (Krasopoulos and Bougoffa, 2022)).
Let $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(n)}$ be $n\geq 2$ positive vectors of $d$ real numbers each, such that $0<m_{i}\leq v^{(i)}_{j}\leq M_{i},i\in[n],j\in[d]$. Then
$\prod_{i=1}^{n}\|\mathbf{v}^{(i)}\|_{2}\leq\frac{{\sqrt{d}}^{n-2}}{2^{n-1}}\left(\prod_{k=2}^{n}B_{k}\right)\|\mathbf{v}^{(1)}\ldots\mathbf{v}^{(n)}\|_{1},$
where $B_{k}=\sqrt{\prod_{i=1}^{k}\frac{M_{i}}{m_{i}}}+\sqrt{\prod_{i=1}^{k}\frac{m_{i}}{M_{i}}}$, for $k\geq 2$.
First, see that for $n=2$, the above yields the Pólya–Szegő inequality, Ineq. (2). Second, note that Ineq. (3) is permutation-sensitive, i.e., it does matter in which order the vectors are input. To see why this is the case, consider the $B_{k}$ constants: they consider only the statistics of the first $k$ vectors.
2.2.1. Improved Version.
To obtain a better lower bound in Ineq. (3), we will consider all (non-redundant) permutations of the input vectors, namely the (n1)!(n-1)! permutations obtained after excluding those obtained by swapping the first two elements, and take the maximum of the resulting lower bounds. While this increases the optimization time, this leads to better bounds. We briefly mention that it is even possible to have a convolution-like bound, by resorting to a subset dynamic programming in time $O(3^{n})$; however, we have not observed any improvement in the bounds.
Figure 3. xBound at runtime: Given a multi-way join on the same key, it (a) lower-bounds the number of joining keys (= $m$), (b) uses $m$ as the prefix length for evaluating degree-sequence norms on the base tables, and (c) takes the best bound from the generalized reverse Hölder’s inequality, min-degree bound, and $m$. The output represents a hard lower bound on the query size.
3. Join Size Lower Bounds
[…]
The degree sequences depicted on the side do not capture the keys for which the frequencies are attained. The question is thus: What is a valid (non-zero) lower bound in this case? A moment of thought reveals that multiplying cross-wise the values in the degree sequences yields the best we can hope for. [^2] See the visualization; this also inspired our estimator’s acronym. In this case, the lower bound is 1 $\times$ 3 + 1 $\times$ 2 + 2 $\times$ 1 = 7. This does have a mathematical argument behind it, namely the rearrangement inequality; we already discussed it in  §2.1.
On the other side, SafeBound (and, extrapolated, LpBound) uses element-wise multiplication to upper bound the join size. In this example, the upper bound reads: 1 $\times$ 1 + 1 $\times$ 2 + 2 $\times$ 3 = 9.
From Degree Sequences to Norms. Surely, we could operate on degree sequences, similar to what SafeBound (Deeds et al., 2023b, a) did for upper bounds in the context of Berge-acyclic queries. However, simply storing them might yield increased storage costs. Instead, we will follow the approach of LpBound, namely operating with the associated norms of the degree sequences. The main disadvantage is that $\ell_{p}$-norms are agnostic to the order of the elements in the vector they are applied to.
Overview. We can visualize xBound at runtime as in Fig. 3: First, (a) we infer a lower bound on the number of distinct joining keys. This is essentially the length of the degree sequences which we should take the norms from. Then, (b) we estimate the norms needed for the reverse inequalities in  §2.2, such as $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$. By estimating, we mean computing a strict lower and upper bound of them. Finally, (c) we lower-bound the dot product by taking the maximum of all bounds output by the set of inequalities. Note that the value from (a) is also a valid lower bound.
Next, we start with the very first step: lower-bounding the number of joining keys.
3.1. Lower-Bounding $\ell_{0}$
Why This is Needed. The attentive reader might have remarked that our toy example in Fig. 2 does not entirely reflect the reality. Indeed, it is not always the case that the two relations will have the same set of keys. As a consequence, if we considered the full key domain, our degree sequences would have been padded with zeros. This is detrimental to the application of the inequalities we presented in  §2.2, as they all require positive vector entries.
Our key insight is that we can first infer a lower bound on the number of joining keys, essentially $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$, and then apply the inequalities with norms confined to the first $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}$ entries in the degree sequences of both relations. Note that this alone is a valid lower bound on the join size. As we will later show, dot-product inequalities can boost it even further.
Two-Way Joins. Let $\mathcal{K}_{A}$ and $\mathcal{K}_{B}$ be the set of keys in $A(X)$ and $B(X)$, respectively. Bounding the number of joining keys is equivalent to bounding $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$. Surely, having at our disposal only the cardinalities of the two sets, $\mathcal{K}_{A}$ and $\mathcal{K}_{A}$, we cannot have a better lower bound than 0. The key realization is that we can use the min/max-values of the columns, commonly referred to as zonemaps (Graefe, 2009), which are a standard feature in modern file formats (3; M. Kuschewski, D. Sauerwein, A. Alhomssi, and V. Leis (2023); 37; A. Afroozeh and P. Boncz (2025)). Let us understand why this changes the whole picture. Note that we can re-write $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$ as follows:
(4) $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|=|\mathcal{K}_{A}|+|\mathcal{K}_{B}|-|\mathcal{K}_{A}\cup\mathcal{K}_{B}|.$
Thus, to lower-bound $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|$ we are to upper-bound $|\mathcal{K}_{A}\cup\mathcal{K}_{B}|$ (due to the negative sign). In turn, the latter can be bounded from above by the span size, i.e., $\max\{\mathcal{K}_{A},\mathcal{K}_{B}\}-\min\{\mathcal{K}_{A},\mathcal{K}_{B}\}$. With this we have:
(5) $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|\geq\max\{0,|\mathcal{K}_{A}|+|\mathcal{K}_{B}|-(\max\{\mathcal{K}_{A},\mathcal{K}_{B}\}-\min\{\mathcal{K}_{A},\mathcal{K}_{B}\})\}.$
Note that real-world systems typically do not store the exact values of $|\mathcal{K}_{A}|$ and $|\mathcal{K}_{B}|$. For example, an optimizer may rely on sketches such as HyperLogLog (Flajolet et al., 2007) to estimate the number of distinct values in a column. In its current form, xBound computes these numbers exactly, separately from the system. We get back to this discussion later, in  §6.
Example. Let us consider an example. Take as domain the set $\{1,\ldots,100\}$, and let the two sets have cardinalities $|\mathcal{K}_{A}|$ = 75 and $|\mathcal{K}_{B}|$ = 50. What is, in this case, a valid lower bound on the intersection size? At a closer inspection, the two sets can be at the two ends of the range, and thus they intersect only on 25 elements. This exactly matches Eq. (5): $|\mathcal{K}_{A}\cap\mathcal{K}_{B}|\geq 75+50-100=25$.
We now outline a simple generalization to multi-way joins.
Multi-Way Joins. Note that we can generalize Eq. (4) as follows. Given $n$ sets $\{\mathcal{K}_{i}\}_{i=1}^{n}$, we are to lower-bound their intersection size. Note again that we do not dispose of the sets themselves, but of limited statistics related to them, such as (bounds on) their cardinalities and their domain size.
To this end, we will be employing a lightweight generalization of the two-way join case. Denote the universe as $U\vcentcolon=\bigcup_{i}^{n}\mathcal{K}_{i}$. Then, it holds:
$\bigl|\;\bigcap_{i=1}^{n}\mathcal{K}_{i}\;\bigr|$ $=|U|-\Bigl|\;\bigcup_{i=1}^{n}\;(U\setminus\mathcal{K}_{i})\;\Bigr|$
$\geq|U|-\sum_{i=1}^{n}|U\setminus\mathcal{K}_{i}|$
$=|U|-\sum_{i=1}^{n}\bigl(|U|-|\mathcal{K}_{i}|\bigr)$
$=|U|-n|U|+\sum_{i=1}^{n}|\mathcal{K}_{i}|$
(6) $=\sum_{i=1}^{n}|\mathcal{K}_{i}|-(n-1)|U|.$
Note that, for the case $n=2$, we obtain exactly the two-way join case, Eq. (4).
This concludes our approach to lower bounding the number of joining keys for an arbitrary number of joins on the same join key. In a subsequent section,  §4.2, we will further develop it, by partitioning the value range into equi-width bins, and applying the above strategy within each bin. Until then, we will first address the aspect of how to leverage this lower bound to enable reverse dot-product inequalities, which will further boost this value.
3.2. Enabling Reverse Inequalities
We show how the last section leads to a valid application of the reverse inequalities in  §2.2, which require that the vectors have positive entries only.
Intuition. Let us start with the simplest case, the two-way join. Let $A(X,\ldots)$ and $B(X,\ldots)$ be two relations which join on $X$, and let $D$ be the support of both $A(X)$ and $B(X)$. By the above considerations (§3.1), we know the lower bound $\ell_{A\,\bowtie\,B,\,X,\,0}^{-}\vcentcolon=m$ on the number of distinct $X$ keys of the join $A\bowtie B$. Note that this tells us a little bit more. Indeed, let $\mathbf{a}$ and $\mathbf{b}$ be the degree vectors of column $X$ in $A$ and $B$, respectively. Recall that the lower bound $m$ simply tells us that we can find a set $\mathcal{K}\subseteq A(X)\cap B(X)$ of size $|K|=m$, such that $\mathbf{a}[k]\cdot\mathbf{b}[k]\neq 0,\forall k\in\mathcal{K}$. This, in turn, implies that both $\mathbf{a}$ and $\mathbf{b}$ take positive values on $\mathcal{K}$, which makes them amenable to dot-product reverse inequalities.
Moreover, the set $\mathcal{K}$ is only a subset of the actual set of joining keys. In other words, the sum $\sum_{k\in K}\mathbf{a}[k]\cdot\mathbf{b}[k]$ is a lower bound on the actual join size, $\sum_{k\in D}\mathbf{a}[k]\cdot\mathbf{b}[k]$.
Two-Way Joins. The above explanation has a caveat: we actually do not have access to the degree vectors themselves. Surely, we could store them, as argued in the beginning of  §3, but this is space-inefficient. What we do have access to are the norms of the (corresponding) degree sequences. With these, we can lower-bound the aforementioned sum $\sum_{k\in K}\mathbf{a}[k]\cdot\mathbf{b}[k]$ by exploiting the prefix of length $m$ in both degree sequences, as follows:
Lemma 3.1 (Prefixes of Degree Sequences; two-way join).
Let $\mathbf{a}$ and $\mathbf{b}$ be the degree vectors of column $X$ in relations $A$ and $B$, respectively, and let $\mathcal{K}$ denote a subset of the joining keys, i.e., $\mathbf{a}[k]\cdot\mathbf{b}[k]\neq 0,\forall k\in\mathcal{K}$. Moreover, let $d^{A}$ and $d^{B}$ be the corresponding degree sequences. Then, there exists a permutation $\sigma:[|\mathcal{K}|]\to[|\mathcal{K}|]$, such that
$\sum_{k\in\mathcal{K}}\mathbf{a}[k]\mathbf{b}[k]\geq\sum_{i=1}^{|\mathcal{K}|}d^{A}_{i}d^{B}_{\sigma_{i}}.$
Proof.
Let $(\hat{a}_{i})_{i=1}^{|\mathcal{K}|}$ and $(\hat{b}_{i})_{i=1}^{|\mathcal{K}|}$ denote the (non-decreasing) degree sequences of $X$ in $A$ and $B$, respectively, yet with values from $\mathbf{a}[\mathcal{K}]$ and $\mathbf{b}[\mathcal{K}]$ only. Intuitively, these are subsequences of the actual degree sequences $d^{A}$ and $d^{B}$.
Observe now that there is a permutation $\sigma:[|\mathcal{K}|]\to[|\mathcal{K}|]$, such that we can re-write the initial dot-product as
$\sum_{k\in\mathcal{K}}\mathbf{a}[k]\mathbf{b}[k]\triangleq\sum_{i=1}^{|K|}\hat{a}_{i}\hat{b}_{\sigma_{i}}.$
In turn, it also holds that $\hat{a}_{i}\geq d^{A}_{i}$, $\forall i\in[|\mathcal{K}|]$; analogously for $\hat{b}$. With this, we obtain
$\sum_{i=1}^{|K|}\hat{a}_{i}\hat{b}_{\sigma_{i}}\geq\sum_{i=1}^{|K|}d^{A}_{i}d^{B}_{\sigma_{i}},$
which concludes the proof. ∎
This lemma allows us to connect a lower bound on the number of joining keys with a dot-product on the degree sequences. And this is a key observation: We can now apply any dot-product reverse inequality, as degree sequences have positive entries only.
Example. Let us consider a concrete example. We would like to apply the Pólya–Szegő inequality, Ineq. (2), and we know a lower bound $m=4$ on the number of joining keys. Consider the two degree sequences below:
$d^{A}=({\color[rgb]{0,0,1}1},{\color[rgb]{0,0,1}1},{\color[rgb]{0,0,1}1},{\color[rgb]{0,0,1}2},{\color[rgb]{0,0,1}2},{\color[rgb]{0,0,1}3}),\qquad d^{B}=({\color[rgb]{1,.5,0}1},{\color[rgb]{1,.5,0}1},{\color[rgb]{1,.5,0}2},{\color[rgb]{1,.5,0}3},{\color[rgb]{1,.5,0}3},{\color[rgb]{1,.5,0}4}).$
The lower bound using the degree sequences reads: 1 $\times$ 3 + 1 $\times$ 2 + 1 $\times$ 1 + 2 $\times$ 1 = 8, as we only consider the prefix of length $m=4$. Recall, however, that we only operate with statistics about these degree sequences. Indeed, to apply the Pólya–Szegő inequality, Ineq. (2), we need the values of $\ell_{2},\ell_{\infty}$, and $\ell_{-\infty}$. Calculating them from the (prefixed) degree sequences, we obtain
$\begin{array}[]{c|c|c|c}\text{degree sequence}&\ell_{2}&\ell_{\infty}&\ell_{-\infty}\\ \hline\cr d^{A}[:4]&\sqrt{7}&2&1\\ d^{B}[:4]&\sqrt{15}&3&1\end{array}$
Therefore, we can lower-bound the join cardinality as
$|A\bowtie B|\geq 2\frac{{\color[rgb]{0,0,1}\sqrt{7}}\cdot{\color[rgb]{1,.5,0}\sqrt{15}}}{\sqrt{\frac{{\color[rgb]{0,0,1}2}\cdot{\color[rgb]{1,.5,0}3}}{{\color[rgb]{0,0,1}1}\cdot{\color[rgb]{1,.5,0}1}}}+\sqrt{\frac{{\color[rgb]{0,0,1}1}\cdot{\color[rgb]{1,.5,0}1}}{{\color[rgb]{0,0,1}2}\cdot{\color[rgb]{1,.5,0}3}}}}\geq 7.17.$
Note that this is slightly worse than the lower bound via the degree sequences directly. This concludes the example for two-way joins.
The topic of actually providing the base table norms will be the topic of  §4. So far, we simply described how to enable dot-product reverse inequalities, by exploiting the lower bound on the number of joining keys calculated in  §3.1. Next, we will generalize the above mechanism to multi-way joins.
Multi-Way Joins. In the sequel, we will be considering a join over the same attribute $X$, i.e., we have $n$ relations $A^{(i)}(X,\ldots)$ and $\mathbf{a}^{(i)}$ the corresponding degree vectors, $i\in[n]$. In addition, denote the domain $D$ as the support of all $A^{(i)}(X)$. Similar to the two-way setting, we will have to first lower bound the number of joining keys that the join produces. This has already been treated in  §3.1. We next show how to improve on this value by using reverse dot-product inequalities.
We will just need a generalization of the two-way case. Let $m$ be the lower bound on its cardinality as in  §3.1, i.e.,
$m\vcentcolon=\ell_{\bowtie_{i=1}^{n}A^{(i)},\>X,\>0}^{-}.$
Recall that this means that we can find a subset $\mathcal{K}$ of the joining keys set, i.e., $\mathcal{K}\subseteq\cap_{i=1}^{n}A^{(i)}$. In turn, the sum $\sum_{k\in\mathcal{K}}\prod_{i=1}^{n}\mathbf{a}^{(i)}[k]$ is indeed a lower bound on the actual join size, $\sum_{k\in D}\prod_{i=1}^{n}\mathbf{a}^{(i)}[k]$, since $\mathcal{K}\subseteq D$. To come again to the realm of degree sequence norms, we need a similar result to Lemma 3.1:
Lemma 3.2 (Prefixes of Degree Sequences; multi-way join).
Let $\mathbf{a}^{(i)}$ and $d^{(i)}$ be the degree vectors and sequences, respectively, of column $X$ in relations $A^{(i)}$, $\forall i\in[n]$, and $\mathcal{K}$ a subset of the joining keys, i.e., $\prod_{i=1}^{n}\mathbf{a}^{(i)}[k]\neq 0,\forall k\in\mathcal{K}$. Then, there exist permutations $\sigma^{(i)}:[|\mathcal{K}|]\to[|\mathcal{K}|]$, $\forall i\in[2,n]$, such that
$\sum_{k\in\mathcal{K}}\prod_{i=1}^{n}\mathbf{a}^{(i)}[k]\geq\sum_{j=1}^{|\mathcal{K}|}d^{(1)}_{j}\prod_{i=2}^{n}d^{(i)}_{\sigma^{(i)}_{j}}$.
Proof.
The proof is a simple generalization of that of Lemma 3.1. ∎
The result tells us that we can move away from degree vectors, and solely work with the prefixes (of length $m$) of the corresponding degree sequences. Similarly to the two-way case, we will be able to apply a reverse dot-product inequality, namely the generalized reverse Hölder’s theorem, Ineq. (3).
3.3. Min-Degree Bound
A natural question is whether the above inequalities are exactly what we need to enable lower bounds. On one hand, the above inequalities have the property of being explanatory, i.e., they will return, as LpBound, an expression which can be debugged. On the other hand, they also require the $\ell_{2}$ norm, which is not available by default on the base tables of a system. Driven by a natural quest for simplicity, we found the following bound, which is reminiscent of the max-degree bound in PANDA (Khamis et al., 2017). We term it min-degree.
Essentially, once we have the lower bound on the number of joining keys, $m$, we can fix a relation $R_{i}$, and assume that all other relations join with $\ell_{-\infty}$ keys each. Formally, this reads as follows:
$\max_{i}\ell_{R_{i},X,1}\displaystyle\prod_{j\neq i}\ell_{R_{j},X,\ell_{-\infty}}.$
Note that norm values are taken, as usual, from the prefix of length $m$ of the corresponding degree sequences.
4. Base Table Statistics
The previous section enabled the application of reverse dot-product inequalities to the setting of degree sequences, and implicitly the norms associated with them. The remaining open question is how to provide these norms to the inequalities. On a high level, given the prefix length $m$, we need to extract the required norms from the degree sequences. In database terminology, we need prefix norms with random access.
Prefixes. The easiest solution is to maintain the norms on prefixes of length a power of two. Indeed, this is resemblant to the prefix trick in LpBound (Zhang et al., 2025, Sec. 5; Optimizations). However, note that this optimization in the context of LpBound is simply a booster for the quality of base table norms it returns; indeed, one could always return the norms of the whole table. For xBound, unfortunately, this is an indispensable component to be able to provide any lower bounds.
An attentive reader would note that simply using the prefix trick does not actually lead to the exact value of the wished-for norms. Surprisingly, we do not need the actual values. Note that we can further lower-bound the RHS of any reverse dot-product inequality by providing hard lower and upper bounds on the requested norms. For instance, to lower bound the RHS side of the Pólya–Szegő inequality, we need (i) lower bounds on $\ell_{2}$, $\ell_{\infty}$, and $\ell_{-\infty}$, and (ii) upper bounds on $\ell_{\infty}$ and $\ell_{-\infty}$. We could infer these lower bounds on the norms by taking a look at the immediate prefixes sizes around $m$, the provided prefixed length. Indeed, for $m$ = 6, we could take as a lower bound for $\ell_{2}$ the value of length 4, while for an upper bound of $\ell_{\infty}$, we could take the value of the prefix length 8, as the degree sequence is sorted in non-decreasing order. While this suffices for correctness, i.e., these are hard lower and upper bounds on the requested norms, there is a way how to further boost their quality. We call this trick norm stitching.
4.1. Norm Stitching
Figure 4. Norm stitching: Deriving $\ell_{2}$ lower bounds for non-power-of-two prefixes by extending the previous power-of-two prefix using its $\ell_{\infty}$ value.
The idea is to “stitch” the norm values at the prefixes of length power of two. Consider Fig. 4, which visualizes the idea behind norm stitching. We would like to extract the norm values for $m=6$. Yet, we only have the values at prefixes 4 and 8. To this end, to obtain a lower bound on $\ell_{2}$, as requested by the Pólya–Szegő inequality, we can take the $\ell_{2}$-value at 4, and artificially add (constant) degree values that are guaranteed to be below the actual ones. In this case, we add the degree value $2$ twice, since that is the maximum degree up until prefix length $4$, and we know that the following degree values will be at least that. We thus obtain: $(\sqrt{10})^{2}+2^{2}\times(6-4)=18$. By taking the root, we obtain a lower bound on the actual $\ell_{2}$ value at prefix 6, which is $\sqrt{28}$.
In pseudo-code, norm stitching for $\ell_{2}$ works as follows:
def NormStitching(norms, rel_name, m):
deg_seq_len = 2**i s.t. 2**i = m
l_2 = norms[rel_name][deg_seq_len][’l_2’]
l_infty = norms[rel_name][deg_seq_len][’l_+infty’]
return math.sqrt(
l_2**2 + (m - deg_seq_len) * l_infty**2
Namely, given the prefix length $m$, we fetch the $\ell_{2}$ value for the largest stored prefix (in our case, only power-of-two prefix lengths) as well as its corresponding $\ell_{\infty}$ value. We then stitch the $\ell_{2}$ norm by adding as many $\ell_{\infty}$ contributions as needed to account for the gap.
Downwards Stitching. Notably, it is also possible to do a “downwards” norm stitching: Instead of padding the degree sequence from the preceding power of two, one can pad it from the subsequent power of two. Namely, to obtain a lower bound on $\ell_{2}$ at prefix 6, one can deduct as many $\ell_{\infty}$’s as the gap. In this particular example, we would obtain: $\sqrt{60}^{2}-4^{2}\times(8-6)$ = 28, and by taking the root we obtain, in this particular case, the actual value of the norm. xBound implements both the upwards and downwards stitching strategies, returning the larger of the two resulting bounds.
We now come to the last component of xBound that enhances the quality of the lower bounds on the number of joining keys.
4.2. Partitioning
Lower bounds on the number of joining keys based only on simple statistics, such as the $\ell_{0}$-value of each column, and the respective min/max values, can be rather loose. One solution is to zoom in on the range value to understand where the joining keys actually occur. To this end, we introduce a lightweight partitioning trick that first splits the value range of each column into a fixed number $B$ of equi-width bins, and collects the same statistics as in  §3.1. Note that partitions for which we collect min/max values can be independent of the actual physical partitions.
Let us outline the benefits this approach. First, collecting $\ell_{0}$-statistics within the $B$ ranges is lightweight (for small $B$’s). Second, this provides us with more fine-grained lower bounds on the number of joining keys, leading to better lower and upper bounds on the norm values (especially when using norm stitching).
4.3. Predicate Types
The above considerations hold for non-filtered base tables. We next show how to still provide lower bounds for the most common predicate types.
4.3.1. Equality Predicates.
Similar to LpBound, we consider the most common values (MCVs) of a given predicate column, and build the exact same norms as discussed previously. Concretely, for an MCV $a$ on column $A$ of relation $R$, we compute the set of $\ell_{p}$ norms needed in the reverse inequalities on top of the degree sequence $\text{deg}_{R}(\ast\>|\>X_{i},A=a)$, for all join columns $X_{i}$ of relation $R$.
At query time, for a predicate $A=v$, if $v$ is an MCV, then we simply take the corresponding norms (boosted with norm stitching, eventually). However, if $v$ is not an MCV, we simply return 0-valued norms.
4.3.2. Range Predicates.
We borrow the hierarchical histogram of LpBound, yet adapt it for lower bounds, namely: Each layer is a histogram whose number of buckets is half the number of buckets of the histogram at the layer below and covers the entire domain range of the predicate column $A$. For a histogram bucket with boundaries $x_{i}$, $y_{i}$], we compute the $\ell_{p}$-norms on $\text{deg}_{R}(\ast\>|\>X_{i},A\in[x_{i},y_{i}])$.
At runtime, to fetch the bounds for a range predicate $A\in[x,y]$, we select (i) for the lower bound, the largest bucket that is contained in $[x,y]$, and (ii) for the upper bound, the smallest bucket that contains $[x,y]$.
4.3.3. Conjunctions & Disjunctions.
Let us start with disjunctions, i.e., pred1 or ... or predp, since they are more intuitive. In LpBound, these are estimated by taking, due to Minkowski’s inequality, the sum of the individual norms. Notably, in xBound, since at least one of the disjuncts should qualify, we can simply take the maximum $\ell_{0}$ in the first step of lower-bounding the number of joining keys, $m$. Once $m$ has been reported, we retrieve the norm vectors for all predicates and compute their coordinate-wise maximum; that is, for each norm we take the largest value among the predicates.
Conjunctions, i.e., pred1 and ... and predp, are slightly more delicate, as they can more easily yield a zero lower bound. Since we have to first lower-bound the number of joining keys, $m$, we can treat the conjuncts as individual predicates, and report their $\ell_{0}$ to the first step of lower-bounding the number of joining keys, $m$. Once we have determined $m$, we will do the same strategy as for disjunctions, as it is guaranteed that, for the given $m$ keys, all individual conjuncts evaluate to true.
We continue with the discussion on the support for other predicate types in §6. We next show how to support the above predicates when partitioning (§4.2) is enabled.
Predicates $\times$ Partitioning. To integrate with the partitioning trick, one has to extend the support for the above predicates, by performing the very same trick as described in §4.2 for each MCV and histogram bucket, respectively. We stress that fact that we store per partition solely the statistics related to the distinct key count and the min/max values of the key. In particular, we do not store any $\ell_{p}$-norms. It is indeed possible to extend xBound to do so—with presumably better lower bounds, yet this adds another layer of norms, making the computation of statistics computationally intensive.
5. Evaluation
In this section, we want to experimentally understand (1) whether xBound can unearth any lower bound violations in current database systems, and (2) to what extent it can alleviate the original optimizer’s cardinality underestimates.
Table 1. Overview of a system’s cardinality bounds violations across all supported join subexpressions of a benchmark.
Benchmark System Underest’ed ¡ xBound ¿ LpBound
JOBlight [Uncaptioned image] 85.3% 14.9% 8.3%
[Uncaptioned image] 79.3% 6.9% 0.4%
[Uncaptioned image] 53.4% 0.0% 6.0%
JOBjoin [Uncaptioned image] 91.4% 0.0% 5.0%
[Uncaptioned image] 93.1% 2.7% 0.2%
[Uncaptioned image] 90.7% 0.6% 4.2%
STATS-CEB [Uncaptioned image] 90.1% 1.1% 4.0%
[Uncaptioned image] 70.3% 0.3% 3.4%
[Uncaptioned image] 67.0% 0.0% 5.1%
MSSalesJoin [Uncaptioned image] 10.8% 3.9% 89.2%
Table 2. Correcting systems’ cardinality estimates via lower bound clipping: estimate = max(estimate, xBound). We report the median Q-error for each number of joins, stopping at the largest join size for which clipping still yields improvement.
JOBlight
System 1 join 2 joins 3 joins [Uncaptioned image] 2.34 29.45 417.50 xBound-ed [Uncaptioned image] $\Downarrow$ 2.05 $\Downarrow$ 28.30 417.50 [Uncaptioned image] 2.00 1.61 2.09 xBound-ed [Uncaptioned image] $\Downarrow$ 1.76 1.61 2.09
JOBjoin
System 1 join 2 joins 3 joins [Uncaptioned image] 1.01 2.14 3.39 xBound-ed [Uncaptioned image] 1.01 2.14 3.39 [Uncaptioned image] 2.40 2.34 2.25 xBound-ed [Uncaptioned image] $\Downarrow$ 1.00 $\Downarrow$ 2.06 $\Downarrow$ 2.06
STATS-CEB
System 1 join 2 joins 3 joins [Uncaptioned image] 6.72 77.23 518.83 xBound-ed [Uncaptioned image] $\Downarrow$ 6.19 77.23 518.83 [Uncaptioned image] 1.46 2.06 5.07 xBound-ed [Uncaptioned image] 1.46 2.06 5.07
In summary, as we will demonstrate in detail below:
  • 1.
    xBound uncovers lower bound violations in the optimizers of DuckDB 1.4, PostgreSQL 18, and Fabric DW. Depending on the benchmark, the number reaches 15%. These findings may issue a warning to the system in question.
  • 2.
    Via clipping with the lower bound, the xBound’ed systems obtain better one-sided Q-errors, thus showing a first evidence that underestimates could be alleviated before they even happen.
5.1. Experimental Setup
Competitors
We use the traditional optimizers of DuckDB 1.4, PostgreSQL 18, and Fabric Data Warehouse (DW). Notably, these optimizers do not come with any guarantee on the estimates. In addition, we also consider the state-of-the-art pessimistic cardinality estimator LpBound (Zhang et al., 2025), which returns join size upper bounds, thereby serving as a “mirrored” comparator to the lower bounds produced by xBound.
Benchmarks
We use conventional cardinality estimation benchmarks and one production workload: JOBlight (Kipf et al., 2019b), JOBjoin (Zhang et al., 2025), STATS-CEB (Han et al., 2021), and Microsoft’s MSSalesJoin. JOBjoin is a recent variation of JOB (Leis et al., 2015) without table predicates for measuring the quality of pessimistic cardinality estimators. [^3] We consider only the self-join-free queries, as in the same reference. MSSalesJoin is a join-only variation of the analytical production workload from Microsoft’s internal revenue reporting platform MSSales (Alotaibi et al., 2024), using a 10%-sample, in total 500 GB, of the real-world dataset. This is a significantly larger real dataset compared to the other benchmarks, while the queries have similar complexity to JOBjoin.
For all these benchmarks, we work with the SQL subexpressions we (currently) support, namely multi-way joins on the same key. Surprisingly, this is exactly the setup in JOBlight, the benchmark used par excellence by learned estimators (Kipf et al., 2019b). In addition, in the case of JOBjoin, we pushed the limit a bit further: given that the subexpressions tend to have a multi-way join on the same join-key which, in turn, joins with unfiltered dimension tables, e.g., info_type, we directly report the size of the respective multi-way join. For an overview of the number of supported join subexpressions and their maximum number of joins, consider Tab. 3.
Table 3. Overview of benchmarks used in the experiments: (i) total number of benchmark queries, (ii) ratio of supported subexpressions, and (iii) the maximum number of joins in the supported subexpressions.
Benchmark #$Q$’s %supported-subexpr’s max. #joins
JOBlight 70 696 / 696 4
JOBjoin 31 9566 / 9566 16
STATS-CEB 146 1830 / 3235 4
MSSalesJoin 47 508 / 50k 3
System configuration
We conduct the experiments on the open systems on a single node equipped with an Intel® Xeon® Gold 5318Y CPU (24 cores, 48 hyper-threads) processor and 128 GB DDR4 main memory, running Ubuntu 24.04. The experiments on Fabric Warehouse (DW) are conducted on Microsoft Fabric with a F256 capacity and version 12.0.2000.8.
Unless otherwise specified, xBound uses the following default configuration: 256 partitions, 5000 MCVs for equality predicates, and 128 histogram buckets for range predicates (the latter two are exactly those from LpBound). Moreover, we use LpBound with $\ell_{p}$ norms $\in\{\ell_{0},\ell_{1},\ell_{2},\ell_{\infty}\}$ to enable a fair comparison by applying the same set of norms to both the lower and upper bounds. [^4] In a strict sense, $\ell_{-\infty}$ is not a norm; see §2. We enable the prefix trick in LpBound, yet disable its PK-FK optimization (Ref. (Zhang et al., 2025, Sec. 5)), i.e., computing norms on two-way joins, since this incurs a high overhead, especially when building LpBound on MSSalesJoin. Both estimators store their norms into the norms tables, while xBound’s partitioned $\ell_{0}$-statistics are stored in a separate table.
Additionally, on MSSalesJoin, the partitioned $\ell_{0}$-statistics are computed in an approximate manner via a HyperLogLog sketch (Flajolet et al., 2007), as exact computation would have been too costly; we discuss the implication of this aspect in §6.
5.2. Mind the Bound
As a warm-up, we want to understand the quality of today’s system estimators, specifically, whether they lie within the safe range defined by the corresponding lower and upper bounds. The experiment is summarized in Tab. 1, where we show the percentage of bound violations per (benchmark, system) pair, when using both xBound’s lower bounds and LpBound’s upper bounds, respectively. Note that we also report the percentage of underestimates. As we see, there is a tendency to mostly underestimate, as already underlined by Leis et al. (Leis et al., 2015). It is precisely this problematic issue that xBound aims to address.
Lower bound violations.
Due to its early stage of development, DuckDB’s estimator is still imprecise. On JOBlight, 14.9% of the estimates lie below the corresponding xBound’s lower bounds. In turn, for PostgreSQL the number reads 6.9%. On the rest of the benchmarks, the numbers look more promising (from the systems’ perspective): On STATS-CEB, xBound finds 1.1% and 0.3% violated lower bounds on DuckDB and PostgreSQL, respectively. On the other side, Fabric DW stands out: Only on JOBjoin could we find a mere of 0.6% of violated lower bounds.
On the other hand, on MSSalesJoin, Fabric DW tends to overestimate (only 10.8% actual underestimates). Out of these 10.8%, only one third violates xBound’s lower bound.
Upper bound violations.
In the same Tab. 1, we also show how the systems perform relative to the corresponding upper bounds. This comparison gives us a first impression whether xBound’s lower bounds are competitive with the state-of-the-art pessimistic cardinality estimator, LpBound. On all research benchmarks, LpBound found upper bound violations across all systems, with PostgreSQL being more conservative. On MSSalesJoin, all Fabric DW’s overestimates are above the corresponding upper bounds.
5.3. Correcting Cardinality Underestimation
Discovering bound violations in estimators is not the sole goal of lower bounds (at least not that of xBound): Next, inspired by the seminal analysis by Leis et al. (Leis et al., 2015), we would like to understand whether having lower bounds can correct the infamous cardinality underestimates. Indeed, they showed more than ten years ago that systems tend to underestimate. Surprisingly, this issue is still present to this very day. Namely, we replicate their evaluation style in Tab. 2, and show the median Q-error for all (supported) subexpressions of a given benchmark, varying the number of joins. Note that, to avoid unnecessary visual clutter, we skip the number of joins from which the estimation errors are the same.
xBound-ed systems
The main observation is that xBound-ed systems, compared to the vanilla ones, have better one-sided Q-errors, i.e., xBound’s lower bounds help them approach the actual cardinality (from below). For instance, when estimating JOBlight’s subexpressions on DuckDB, the median Q-error for one join goes down from 2.34 to 2.05. On PostgreSQL, on the same number of joins, the error drops from 2.00 to 1.76. On STATS-CEB, xBound could improve DuckDB’s median Q-error from 6.72 to 6.19. Notably, on JOBjoin, xBound still manages to lower the error for three joins, namely from 2.25 to 2.06. On the rest of the benchmarks, given the limited number of lower bound violations found, we could not see any Q-error improvement.
5.4. Statistics Overhead
Relying on $\ell_{p}$-norms and min/max-values makes xBound a lightweight estimator, similar to LpBound. Indeed, in Tab. 4, we outline the overhead to build and store xBound statistics on a given DuckDB database. In terms of the database itself, we report the time needed to load it from the benchmark’s CSV files, and the resulting database file. We also report LpBound’s norms table size, and the necessary time to build the statistics.
Table 4. Space and build overhead of xBound (256 partitions).
Benchmark Database xBound LpBound
norms part. $\ell_{0}$
Size Load Size Build Size Build Size Build
[MB] [s] [MB] [s] [MB] [s] [MB] [s]
JOBlight 2646.4 50.5 5.8 26.8 21.2 27.9 2.6 6.0
JOBjoin 2646.4 50.5 0.8 24.9 ¡0.1 5.4 ¡0.1 6.2
STATS-CEB 22.9 1.2 9.4 14.3 34.3 247.5 7.6 7.8
We observe that the sizes of the norms tables of both estimators are on par. Indeed, xBound has the overhead of the prefix trick, since, while LpBound does it only for the non-filtered base tables, we have to do it in xBound even for MCV and histograms, so as to provide non-zero lower bounds. However, this extra layer of prefixes can be observed in the build time. It is worth mentioning that since xBound and LpBound share the same MCV and histogram tables, the overhead of computing them can be amortized when they are built together. In particular, because both estimators operate on the very same degree sequence (albeit sorted in different orders), one could merge the two statistics calculators to accelerate the build procedure.
Unlike LpBound, xBound also has to build, in case partitioning (§4.2) is enabled, the partitioned $\ell_{0}$-statistics. As we strive for exact distinct key counts, this will also incur an overhead, yet still comparable to that of building the actual norms. Note that the STATS-CEB benchmark has a larger number of range predicates, hence the increased overhead.
Figure 5. Average estimation time of systems and estimators for varying join sizes on all subexpressions of JOBlight.
5.5. Estimation Time
By relying on a handful of statistics, xBound’s estimation is comparable to that of traditional estimators, such as those used in DuckDB and PostgreSQL. To understand the dependency of the estimation time with respect to the number of joins, consider Fig. 5, where we plot the optimization time of the two systems, side by side with that of LpBound ($p$ = 2) and xBound. For both systems, we use the time spent to produce the plan (via EXPLAIN PLAN). For LpBound, we report only the solving time of the linear program, whereas for xBound, we measure the time taken by its two components: (i) estimating the distinct key count of the join, and (ii) applying the reverse dot-product inequalities. Note that since the lower bounds in step (i) tend to be 0 starting with 4 joins, the estimation times remain constant.
5.6. Downstream Benefits of xBound
Having shown that xBound’s lower bounds can exceed Fabric DW’s cardinality estimates, we now quantify the downstream benefit to reduce spilling—writing internal operator state to disk due to underestimating their memory consumption as consequence of cardinality underestimation. We focus on the MSSalesJoin production workload due to its realistic table/query sizes. We use Fabric DW’s query optimizer test setup to estimate operator memory consumption under the original estimates and under xBound-capped estimates (in this case, xBound uses 1024 partitions; see the ablation study in §5.7), as a function of the projected columns.
In Fig. 6, we report the resulting spilling reduction as a function of how strongly the consuming operators’ state sizes depend on the intermediate’s projected columns: from 0.3 GB when essentially none of the projected input columns contribute to state size, up to 94 GB when all columns do. The trend is linear, since the memory footprint scales with average row size times cardinality, and increasing the fraction of projected columns increases row size proportionally. Overall, even modest caps on intermediate cardinalities can translate into substantial reductions in spilling, with the realized benefit governed by the intermediate’s row width.
Figure 6. End-to-end evaluation in Fabric DW: Expected spilling reduction for MSSalesJoin when capping subexpression estimates with xBound’s lower bounds. Results vary with operator state size, which depends on how many input columns are projected.
5.7. Ablation Study
We would like to understand the effect partitioning (§4.2) has on the quality of the bounds, along with the additional optimization of norm stitching (§4.1). In the following, we took MSSalesJoin as our workload, so as to test the effect at large scale. To this end, we varied the number of partitions $\in\{1,4,16,256,1024\}$ and disabled norm stitching.
We show the results of the experiment in Fig. 7, where we plot the Q-error of xBound. [^5] We use boxen plots (Hofmann et al., 2017) since they better expose the full error distribution. The first observation is that number of partitions greatly impacts the quality of the bounds. This is due to fact that, by range partitioning, the joining keys can be better localized, hence leading to better lower bounds in xBound’s first step (§3.1). In addition, norm stitching, while orthogonal to the partitioning trick, can further boost the bounds, especially when the number of partitions is small.
Figure 7. Analyzing the impact of partitioning (§4.2) and norm stitching (§4.1) on the quality of the lower bounds on MSSalesJoin.
6. Discussion
We have demonstrated that cardinality lower bounds hold the potential to solve the long-standing cardinality underestimation problem, still present in current systems despite extensive research. In the following, we will discuss several avenues for generalizing and refining these lower bounds.
Query Type. Currently, we support only Berge-acyclic queries, with the additional requirement that the joins happen on the same join key. While this condition inherently holds for (single-key) two-way joins, e.g., the very first join in a the query plan, and, surprisingly, for the well-studied JOBlight benchmark (Kipf et al., 2019b), real-world queries often involve joins on multiple join keys.
Moreover, while rather common in established cardinality estimation benchmarks, like JOB (Leis et al., 2015) and STATS-CEB (Han et al., 2021), Berge-acyclicity no longer holds in enterprise settings to such a large extent. For instance, many two-way joins in MSSales are, in fact, not Berge-acyclic (yet, still acyclic). It is interesting to understand whether this framework can also be extended to handle cyclic queries, such as those in the Subgraph-Matching benchmark (Sun and Luo, 2020).
Another intriguing extension is that to nested subqueries. While this remains an open question for LpBound, the reverse inequalities used by xBound can be extended to return lower bounds on more norms beyond $\ell_{1}$, in particular $\ell_{2}$, e.g., by using the $\ell_{4}$-norms of the input vectors. In turn, these bounds can be propagated further to subsequent subqueries.
Beyond Inner Joins. To be applicable to an even more diverse workload, xBound should also support non-inner joins. We believe that the framework is flexible enough to be extended in this direction. For example, we can easily support semi-join operators, $A\ltimes B$: the lower bound on the number of distinct keys of the join $A\Join B$3.1) can be used as the prefix length when fetching the $\ell_{1}$ norm of $A$. Moreover, for outer joins, one can always return the lower bound of the (implicit) inner join. However, achieving tighter lower bounds would require more information about the number of non-joining keys, which we leave as future work.
Another relevant operator is group-by. While methods exist to provide approximate estimates either via sketches (Freitag and Neumann, 2019) or learning (Kipf et al., 2019a), and even to provide upper bounds on its cardinality (Zhang et al., 2025), we are not aware of any non-probabilistic lower bounds on the cardinality of this operator.
Predicate Support. While we support the standard predicate types (equality, range, conjunctions, and disjunctions), providing (good) lower bounds for string predicates, such as LIKE and REGEX, remains an open challenge. A simple technique is, naturally, sampling. The number of tuples that qualified within a sample is, indeed, a (loose) lower bound (provided that we do not extrapolate this number to the full dataset).
To generate the actual norms needed in the joins, we have to lower-bound the distinct count of the joining key (similar to the way how MCVs work, §4.3.1). The question is whether one can do better than this. We note that cardinality upper bounds can be derived using tri-grams, as hinted in SafeBound and LpBound, but applying similar techniques for lower bounds remains an open problem.
Theory Advancements. As an attentive reader might have already observed, we are using a rather limited number of statistics: standard $\ell_{p}$-norms, i.e., $\{\ell_{0},\ell_{1},\ell_{2},\ell_{\infty}\}$, along with $\ell_{-\infty}$, and (partitioned) min/max values. This limited set is dictated by the class of inequalities that we are currently using. From a theory point of view, we believe that it is possible to design dot-product reverse inequalities that consider a wider range of norms, as LpBound does. Indeed, this is orthogonal to join size estimation, yet any new inequality could be easily plugged in xBound (if the corresponding norms have been computed).
Deteriorating Lower Bounds. Our lower bounds rely on min/max statistics, or the so-called zonemaps (Graefe, 2009), which are readily available in database catalogs and modern file formats. While simplicity is good for real systems, collecting more granular statistics about the value distribution (not of the degree distribution, already captured by $\ell_{p}$-norms), could lead to better bounds. Indeed, their use in xBound is tightly linked to the quality of the lower bounds on the number of distinct keys in the join. Ultimately, the goal is to “locate” the joining keys more precisely, thereby mitigating the problem of deteriorating lower bounds as the number of joins increases.
xBound+HLL. To test xBound on the MSSales enterprise workload, we had to resort to approximation. Namely, instead of computing the actual numbers, we approximated the distinct key count in the partitioned statistics with HyperLogLog (Flajolet et al., 2007). While the build phase becomes faster than computing the exact distinct counts, this compromises the strict theoretical guarantees. The resulting “lower bounds” hold only in an approximate sense. We would like to understand the guarantees of xBound+HLL in a future work: Can we quantify the error of these approximate lower bounds?
Data Updates. Lower bounds have an intriguing property when it comes to updates. Namely, inserts (and appends) are supported by default, since a lower bound on $|A\Join B|$ also holds when $\Delta_{A}$ is appended to $A$. This contrasts with the struggle of upper bound estimators, such as LpBound, where new data renders the previous degree norms immediately stale. Naturally, the challenge is how to incrementally maintain lower bounds on changing data. One key observation is the following: $|(A+\Delta_{A})\Join B|=|A\Join B|+|\Delta_{A}\Join B|$, due to the fact that the join size is a dot product between two vectors. Consequently, we can take the previous lower bound $|A\Join B|^{-}$, and only add the one for the newly-inserted data, $|\Delta_{A}\Join B|^{-}$. In practice, the deltas should be periodically merged to prevent storage and runtime overhead.
7. Related Work
While, to the best of our knowledge, xBound is the first estimator for nontrivial lower bounds on join result sizes, upper bounds on join sizes have been extensively studied; we summarize the most relevant work at the end of this section. It is also important to distinguish our goal from the literature on fine-grained lower bounds for conjunctive queries (Mengel, 2025; Fan et al., 2023, 2024). Those results show that every algorithm evaluating a given class of conjunctive queries must incur at least a certain complexity under specific computational assumptions. In other words, they yield asymptotic lower bounds, whereas our focus is on data-dependent lower bounds on join cardinalities.
Learned Estimators. Apart from the provable estimators in the last two decades, there has been a long line of research on learned cardinality estimation (Kipf et al., 2019b; Yang et al., 2019; Hilprecht et al., 2020; Wu et al., 2023); see, for instance, the recent analysis by Heinrich et al. (Heinrich et al., 2025). However, these estimators do not come with guarantees on their hard estimates, and can lead to arbitrarily mis-estimations when faced with complex workloads (Wehrstein et al., ). An exception is FactorJoin (Wu et al., 2023), which, indeed, provides probabilistic upper bounds.
Sketch-Based Estimators. Seeing the join size as a dot-product between the two corresponding degree vectors is not something new: Sketch-based estimators (Ganguly et al., 2004, 2005; Wang et al., 2023; Heddes et al., 2024) already use this very observation. Yet, they rely on well-established sketches, which cannot guarantee lower bounds on the actual join size; they solely return an approximation, which could, itself, violate the bounds.
Robust Query Processing. The impossibility to solve the cardinality estimation problem, let our community transition to a new paradigm, where the cardinality estimator is considered to be of secondary importance. This line of research started with Yannakakis’ algorithm (Yannakakis, 1981), and has seen a revival with a series of Dagstuhl seminars (Graefe et al., 2010), and more recently, with concrete implementations in real systems (Yang et al., 2024; Birler et al., 2024; Wang et al., 2025; Zhao et al., 2025; Bekkers et al., 2025).
Join Size Upper Bounds. Recall that an upper bound for a conjunctive query $Q$ is a numerical value, which is computed in terms of statistics on the input database, and is guaranteed to be greater or equal to the output size of any given query. The upper bound is tight if there is a database instance, satisfying the statistics, such that the query’s output size is as large as the bound.
To this end, the AGM bound (Atserias et al., 2008) is a tight upper bound that uses only the cardinalities of the relations; in other words, it uses only the $\ell_{1}$-norms of full degree sequences. To upper-bound the output size, the AGM bound uses fractional edges covers and is useful for cyclic queries.
With PANDA, Khamis, Ngo and Suciu generalize the AGM bound by also employing the $\ell_{\infty}$-norm (Khamis et al., 2017). When applied to acyclic queries, their max-degree bound is an improvement over the AGM bound, yet, usually, is less accurate than a density-based estimator. For instance, to upper-bound the size of $|A\Join_{X=Y}B|$, max-degree uses the following three inequalities (note that the first one is coming from the AGM bound):
$|A\Join_{X=Y}B|$ $\leq|A|\cdot|B|,$
$|A\Join_{X=Y}B|$ $\leq\|\text{deg}_{A}(\ast\>|\>X)\|_{\infty}\cdot|B|,$
$|A\Join_{X=Y}B|$ $\leq|A|\cdot\|\text{deg}_{B}(\ast\>|\>Y)\|_{\infty}.$
Our min-degree estimator §3.3 takes inspiration from this observation for join size lower bounds.
SafeBound (Deeds et al., 2023b, a) uses simple, full degree sequences and computes a tight upper bound of a Berge-acyclic, full conjunctive query. For example, if $\deg_{A}(\ast\>|\>X)=(a_{1}\geq a_{2}\geq\ldots)$ and $\deg_{B}(\ast\>|\>Y)=(b_{1}\geq b_{2}\geq\ldots)$, then SafeBound returns the following upper bound on the two-way join from above: $|A\Join_{X=Y}B|\leq\sum_{i}a_{i}b_{i}$. Indeed, SafeBound can return a much better bound than the ones above. However, its bounds are not describable by a closed-form formula; it is only given by an algorithm.
The above limitations of SafeBound are essentially addressed by LpBound (Zhang et al., 2025), which, instead of relying on degree sequences, it employs their $\ell_{p}$-norms. First, this eliminates the need to store degree sequences. Second, by solving a linear program, this returns a closed-form formula that can be easily understood.
8. Conclusion
In this work, we set out to address an issue that was identified more than ten years ago by Leis et al. (Leis et al., 2015): Systems tend to underestimate. And they still do so today. In these ten years, the theory community has brought us exciting new results on join size upper bounds. From the perspective of a real system, replacing one’s own estimator with a new one requires considerable courage, particularly when these estimators still exhibit significant one-sided Q-errors. For this reason, upper bounds are best suited for correcting the original optimizer’s overestimates.
Observing this need, we introduced xBound, the very first framework for provable join size lower bounds. While limited to multi-way joins on the same join key, xBound’s lower bounds are practical enough to both unearth bound violations in modern optimizers, as well as improve, for the first time, their one-sided Q-errors (by capping the estimate with the corresponding lower bound). This could be seen even on a production workload, on which xBound fixed 36.1% of Fabric Data Warehouse’s underestimates. Given these observations, we outlined several open points that are worth exploring from both a research and systems perspective.
Whether our systems’ optimizers will completely overcome their Achilles heel in the decade ahead, only time will tell.